# Algorithms in $4$-manifold topology

### v1

@@ 8.2. Algorithms for stable diffeomorphism
We recall the basics of modified surgery in Section 8.1. The main result of this theory (for our interests) is Theorem 8.7, which states that the set of stable diffeomorphism classes of $4$-manifolds with a fixed 1-type $\xi$ (Definition 8.6) is in one-to-one correspondence with $\Omega^{\xi}_{4}/\operatorname{Out}(\xi)$, a quotient of a certain bordism group. If either of the assumptions of Theorems 8.3 or 8.4 is satisfied, we will prove that the following map given by the signature $\sigma$, and the primary invariant $\mathfrak{pri}$ (given by evaluating the fundamental class)
@@ 8.3. Combinatorial modified surgery
  1. (1)
    Type I ((((totally non-spin)))). If $w_{2}(\widetilde{X})\neq 0$, the normal 1-type of $X$ is equivalent to
@@
  2. (2)
    Type II ((((spin)))). If $w_{2}(X)=0$, the normal $1$-type is equivalent to
@@
  3. (3)
    Type III ((((almost spin)))). If $w_{2}(X)\neq 0$, but $w_{2}(\widetilde{X})=0$, there is an element $w_{2}^{\pi}\in H^{2}(B\pi;\mathbb{Z}/2)$, such that $w_{2}(X)$ is pulled back from $w_{2}^{\pi}$ along a 2-connected map $X\to B\pi$. The normal $1$-type $\xi$ is equivalent to the map given by the homotopy pullback
    ${B}$${B\pi}$${\operatorname{BSO}}$${K(\mathbb{Z}/2,2)}$$\scriptstyle{\xi}$
@@
    where $w_{1}$ and $w_{2}$ denote the universal maps for the first and second Stiefel–Whitney classes, respectively. A $1$-smoothing is given by an isomorphism $\pi_{1}(X)\to\pi$ such that $w_{2}(X)$ pulls back from $w_{2}^{\pi}$.∎
@@
[…]
  1. (1)
    $w_{2}^{\pi}=\infty$ if $w_{2}(\widetilde{X})\neq 0$,
  2. (2)
    $w_{2}^{\pi}=0$ if $w_{2}(X)=0$,
  3. (3)
    $w_{2}^{\pi}\neq 0,\infty$ if there exists a $\pi_{1}$-isomorphism $c\colon X\to B\pi^{(5)}$ such that $w_{2}(X)=c^{*}(w_{2}^{\pi})$.
@@
Let $X$ be a triangulated $4$-dimensional manifold, let $\pi$ be a finite group and let $f\colon\pi_{1}(X)\to\pi$ be a homomorphism. Then there is a simplicial map between the two-fold barycentric subdivision of $X$ and the two-fold barycentric subdivision of the geometric realisation of the bar construction $B\pi$, which, on the level of the fundamental groups, agrees with $f$. Further, we can factor this simplicial map as a product of simplicial maps $X\to B\pi_{1}(X)\to B\pi$ such that the first map induces the identity map and the second map induces $f$ on the level of fundamental groups.
@@
  • On the 1-skeleton of $X$, choose a maximal tree $T\subset X_{(1)}$ and send $T$ to the only 00-simplex in $B\pi$. For any other 1-simplex $e$ of $X$, pick any based loop in $T\cup e$ which goes through $e$ exactly once (in the correct direction). This determines an element $g\in\pi_{1}(X)$. We send $e$ to the 1-simplex $|f(g)|$ of $B\pi$.
@@ 8.5. Algorithms for combinatorial modified surgery
In the following we discuss the case $R=\mathbb{Z}$. The case $R=\mathbb{Z}/2$ is dealt with in almost the same way. First, using elementary linear algebra we determine a basis for $H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}$ and for $H^{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}$. The fundamental class of $X$ is given by the sum of all oriented 4-dimensional simplices of $\mathcal{K}$, and capping with it induces the Poincaré duality isomorphism $H^{2}(\mathcal{K};\mathbb{Z})\to H_{2}(\mathcal{K};\mathbb{Z})$ which descends to an isomorphism $H^{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}\to H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}$. The inverse to this isomorphism can be computed, e.g. by computing the isomorphism with respect to the above bases, and then computing the inverse matrix. Use this Poincaré duality inverse map and the standard cup product formula (which is known to descend to cohomology) to obtain a map
$H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}\otimes H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}\to H^{4}(\mathcal{K};\mathbb{Z}).$
@@
An alternative way to compute the signature is to utilise the combinatorial formula given by Ranicki–Sullivan [RS76]. From $\mathcal{K}$ we can construct simplicial chain and cochain groups along with boundary and coboundary operators
@@
  • takes as an input an oriented triangulation $\mathcal{K}$ of a closed, oriented, smooth $4$-manifold $X$ along with data for some isomorphism $\pi_{1}(X)\cong\pi$ ((((see Definition 8.9)))), and
@@
Next, we recall that the Wu formula says that in an oriented $4$-manifold $Y$, for every $y\in H^{2}(Y;\mathbb{Z}/2)$, we have $\langle w_{2}(Y),y\rangle=\lambda^{\mathbb{Z}/2}(y,y)$, see for example [GS99, Exercise 5.7.3]. From the universal coefficient theorem with $\mathbb{Z}/2$ coefficients it follows that this formula determines the Stiefel–Whitney class $w_{2}(Y)$. Thus we can compute $w_{2}({\mathcal{K}})$ and $w_{2}(\widetilde{\mathcal{K}})$ by using Lemma 8.19 to compute the parity of the $\mathbb{Z}_{2}$-intersection form of $\mathcal{K}$ and $\widetilde{\mathcal{K}}$.
@@
${H^{2}(B\pi;\mathbb{Z}/2)}$${H^{2}(X;\mathbb{Z}/2)}$${H^{2}(\widetilde{X};\mathbb{Z}/2)}$00{0}
@@
  • takes as input two oriented triangulations $\mathcal{K}$ and $\mathcal{L}$ of two closed, oriented, smooth $4$-manifolds $M$ and $N$ respectively with the same 1-type data, along with both of their reduced smoothing data and isomorphisms $\pi_{1}(M)\cong\pi$ and $\pi_{1}(N)\cong\pi$ ((((Definition 8.9)))), and
@@
For each $f\in\operatorname{Isom}(\pi_{1}(\mathcal{L}),\pi_{1}(\mathcal{K}))$ we construct a simplicial map $Bf\colon B\pi_{1}(\mathcal{L})\to B\pi_{1}(\mathcal{K})^{(5)}$ (see Lemma 8.12). This gives us a map $(Bf)\circ\iota_{\mathcal{L}}\colon\mathcal{L}\to B\pi_{1}(\mathcal{K})^{(5)}$. Similarly, we make a simplicial map $\iota_{\mathcal{K}}\colon\mathcal{K}\to B\pi_{1}(\mathcal{K})^{(5)}$ which induces the identity on the fundamental groups.
@@
Does $\text{Does }Ax=\alpha\text{ for some }x\in C_{5}^{\operatorname{simp}}(B\pi_{1}(\mathcal{K}))?$
If such an $x\in C_{5}^{\operatorname{simp}}(B\pi_{1}(\mathcal{K}))$ exists we return ‘yes’, otherwise we move on to the next isomorphism. If all isomorphisms have been exhausted, we return ‘no’. ∎
@@
In the case of (1) start by generating a certificate (Definition 8.9) that the fundamental groups are both infinite cyclic by applying Corollary 8.17. Similarly, in the case of (2) start by generating a certificate that the fundamental groups of $X_{1}$ and $X_{2}$ are isomorphic to some group $\pi$ by applying Corollary 8.15. If the fundamental groups are not isomorphic then our manifolds are not stably diffeomorphic.
@@
is an injection. Then we finish the proof by applying the algorithms in Lemma 8.23, and Lemma 8.20 or, in the case of $\pi$ infinite cyclic, only the Algorithm 8.20.
The case of $\pi$ infinite cyclic is handled by using the Atiyah–Hirzebruch spectral sequence for types (I) and (II). Since $H^{2}(B\mathbb{Z};\mathbb{Z}/2)=0$, type (III) is impossible.
@@
[…]
$\Omega_{4}^{\xi}$ $\cong 8\mathbb{Z}\oplus H_{4}(B\pi)\quad\text{for type (III) and $\pi$ of even order},$
$\Omega_{4}^{\xi}$ $\cong 16\mathbb{Z}\oplus H_{4}(B\pi)\quad\text{for type (III) and$\pi$ of odd order}.\qed$

### v2

@@ 8.2. Algorithms for stable diffeomorphism
We recall the basics of modified surgery in Section 8.1. The main result of this theory (for our interests) is Theorem 8.7, which states that the set of stable diffeomorphism classes of $4$-manifolds with a fixed 1-type $\xi$ (Definition 8.6) is in one-to-one correspondence with $\Omega^{\xi}_{4}/\operatorname{Out}(\xi)$, a quotient of a certain bordism group. If either of the assumptions of Theorems 8.3 and 8.4 is satisfied, we will prove that the following map given by the signature $\sigma$, and the primary invariant $\mathfrak{pri}$ (given by evaluating the fundamental class)
@@ 8.3. Combinatorial modified surgery
  • Type I
    (totally non-spin). If $w_{2}(\widetilde{X})\neq 0$, the normal 1-type of $X$ is equivalent to
@@
  • Type II
    (spin). If $w_{2}(X)=0$, the normal $1$-type is equivalent to
@@
  • Type III
    (almost spin). If $w_{2}(X)\neq 0$, but $w_{2}(\widetilde{X})=0$, there is an element $w_{2}^{\pi}\neq 0\in H^{2}(B\pi;\mathbb{Z}/2)$, such that $w_{2}(X)$ is pulled back from $w_{2}^{\pi}$ along a 2-connected map $X\to B\pi$. The normal $1$-type $\xi$ is equivalent to the map given by the homotopy pullback
[…]
@@
    where $w_{2}$ denotes the universal map for the second Stiefel–Whitney class. A $1$-smoothing is given by an isomorphism $\pi_{1}(X)\to\pi$ such that $w_{2}(X)$ pulls back from $w_{2}^{\pi}$.∎
@@
where $T_{\pi}$ is a multiplication table for the finite group $\pi$, $(B\pi)^{(5)}$ is the 5-skeleton of some triangulation of some $B\pi$, and $w_{2}^{\pi}$ is an element of the set $H^{2}(B\pi^{(5)};\mathbb{Z}/2)\cup\{\infty\}$, where $\{\infty\}$ is a set with one element disjoint from the cohomology group $H^{2}(B\pi^{(5)};\mathbb{Z}/2)$. We require that $w_{2}^{\pi}=\infty$ if $w_{2}(\widetilde{X})\neq 0$, while if $w_{2}(\widetilde{X})=0$, we require that there exists a $\pi_{1}$-isomorphism $c\colon X\to B\pi^{(5)}$ such that $c^{*}(w_{2}^{\pi})=w_{2}(X)\in H^{2}(X;\mathbb{Z}/2)$.
@@
Let $X$ be a triangulated $4$-dimensional manifold, let $\pi$ be a finite group and let $f\colon\pi_{1}(X)\to\pi$ be a homomorphism. Then there is a simplicial map between the two-fold barycentric subdivision of $X$ and the two-fold barycentric subdivision of the geometric realisation of the bar construction $B\pi$, which, on the level of fundamental groups, agrees with $f$. Further, we can factor this simplicial map as a composition of simplicial maps $X\to B\pi_{1}(X)\to B\pi$ such that the first map induces the identity map and the second map induces $f$ on the level of fundamental groups.
@@
  • On the 1-skeleton of $X$, choose a maximal tree $T\subset X_{(1)}$ and send $T$ to the only 0-simplex in $B\pi$. For any other 1-simplex $e$ of $X$, pick any based loop in $T\cup e$ which goes through $e$ exactly once (in the correct direction). This determines an element $g\in\pi_{1}(X)$. We send $e$ to the 1-simplex $|f(g)|$ of $B\pi$.
@@ 8.5. Algorithms for combinatorial modified surgery
In the following we discuss the case $R=\mathbb{Z}$. The case $R=\mathbb{Z}/2$ is dealt with in almost the same way. First, from $\mathcal{K}$ we construct the simplicial chain and cochain groups along with boundary and coboundary operators. Using this data, we then compute the homology and cohomology groups. Note that cup and cap products are defined on, and can be computed on, the (co)chain level via standard formulae, and hence we can use these operations in our algorithm. Using elementary linear algebra we determine a basis for $H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}$ and for $H^{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}$. The fundamental class of $X$ is given by the sum of all oriented 4-dimensional simplices of $\mathcal{K}$, and capping with it induces the Poincaré duality isomorphism $H^{2}(\mathcal{K};\mathbb{Z})\to H_{2}(\mathcal{K};\mathbb{Z})$ which descends to an isomorphism $H^{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}\to H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}$. The inverse to this isomorphism can be computed, e.g. by computing the isomorphism with respect to the above bases, and then computing the inverse matrix. Use this Poincaré duality inverse map and the standard cup product formula (which is known to descend to cohomology) to obtain a map
$H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}\times H_{2}(\mathcal{K};\mathbb{Z})/\operatorname{Tors}\to H^{4}(\mathcal{K};\mathbb{Z}).$
@@
An alternative way to compute the signature is to utilise the combinatorial formula given by Ranicki–Sullivan [RS76]. As in the proof of Lemma 8.19 we can construct simplicial chain and cochain groups along with boundary and coboundary operators
@@
  • takes as an input an oriented triangulation $\mathcal{K}$ of a closed, oriented, smooth $4$-manifold $X$ along with data for some isomorphism $\pi_{1}(X)\cong\pi$ ((see Definition 8.9)), and
@@
[…]
@@
$H^{2}(B\pi;\mathbb{Z}/2)$$H^{2}(X;\mathbb{Z}/2)$$H^{2}(\widetilde{X};\mathbb{Z}/2)$}0{0}
@@
  • takes as input two oriented triangulations $\mathcal{K}$ and $\mathcal{L}$ of two closed, oriented, smooth $4$-manifolds $M$ and $N$ respectively with the same 1-type data, along with both of their reduced smoothing data and isomorphisms $\pi_{1}(M)\cong\pi$ and $\pi_{1}(N)\cong\pi$ ((Definition 8.9)), and
@@
For each $f\in\operatorname{Isom}(\pi_{1}(\mathcal{L}),\pi_{1}(\mathcal{K}))$ we construct a simplicial map $Bf\colon B\pi_{1}(\mathcal{L})^{(5)}\to B\pi_{1}(\mathcal{K})^{(5)}$ (see Lemma 8.12). This gives us a map $(Bf)\circ\iota_{\mathcal{L}}\colon\mathcal{L}\to B\pi_{1}(\mathcal{K})^{(5)}$. Similarly, we make a simplicial map $\iota_{\mathcal{K}}\colon\mathcal{K}\to B\pi_{1}(\mathcal{K})^{(5)}$ which induces the identity on the fundamental groups.
@@
$\text{Does }Ax=\alpha\text{ for some }x\in C_{5}^{\operatorname{simp}}(B\pi_{1}(\mathcal{K})^{(5)})?$
If such an $x\in C_{5}^{\operatorname{simp}}(B\pi_{1}(\mathcal{K})^{(5)})$ exists we return ‘yes’, otherwise we move on to the next isomorphism. If all isomorphisms have been exhausted, we return ‘no’. ∎
@@
In the case of (1), start by generating a certificate (Definition 8.9) that the fundamental groups are both infinite cyclic by applying Corollary 8.17. Similarly, in the case of (2), start by generating a certificate that the fundamental groups of $X_{1}$ and $X_{2}$ are isomorphic to some group $\pi$ by applying Corollary 8.15. If the fundamental groups are not isomorphic then our manifolds are not stably diffeomorphic.
@@
is an injection. Then we finish the proof by applying the algorithms in Lemmas 8.23 and 8.20 or, in the case of $\pi$ infinite cyclic, only Algorithm 8.20.
The case (1) of $\pi$ infinite cyclic is handled by using the Atiyah–Hirzebruch spectral sequence for types (I) and (II). Since $H^{2}(B\mathbb{Z};\mathbb{Z}/2)=0$, type (III) is impossible.